# SAKE: Steering Activations for Knowledge Editing

### v1

@@ Abstract
As Large Langue Models have been shown to memorize real-world facts, the need to update this knowledge in a controlled and efficient manner arises. Designed with these constraints in mind, Knowledge Editing (KE) approaches propose to alter specific facts in pretrained models. However, they have been shown to suffer from several limitations, including their lack of contextual robustness and their failure to generalize to logical implications related to the fact. To overcome these issues, we propose SAKE, a steering activation method that models a fact to be edited as a distribution rather than a single prompt. Leveraging Optimal Transport, SAKE alters the LLM behavior over a whole fact-related distribution, defined as paraphrases and logical implications. Several numerical experiments demonstrate the effectiveness of this method: SAKE is thus able to perform more robust edits than its existing counterparts.
@@
Marco Scialanga*,1,2, Thibault Laugel*,1,3, Vincent Grari1,3, Marcin Detyniecki1,3,4 1AXA, Paris, France, 2EPFL, Lausanne, Switzerland, 3TRAIL, LIP6, Sorbonne Université, Paris, France 4Polish Academy of Science, IBS PAN, Warsaw, Poland *Equal contribution Correspondence: marco.scialanga@epfl.ch, thibault.laugel@axa.com
@@ 2.1 Knowledge Editing
Knowledge Editing (see Wang et al. (2024a) for a survey) is the task of modifying the knowledge stored in LLMs. Given a language model $f$, an edit is generally defined by a source tuple $(s,r,o)$ ($s$ being the subject, $r$ the relation, and $o$ an old object), and a target tuple $(s,r,o^{*})$, with $o^{*}$ a new object. Concretely, the object $o$ is to be returned when querying the model with $(s,r)$ in the form of natural language (the prompt $p$): the goal of model editing is thus to define an edited model $f^{*}$ that outputs $o^{*}$ given the prompt $p$, an objective generally defined as "edit success" or "accuracy". Besides accuracy, two other objectives are generally associated to the task of knowledge editing Meng et al. (2022a); Wang et al. (2024a): the ability of the model to generate $o^{*}$ given paraphrases $p^{\prime}$ of $p$ (generality), and how well the initial behavior of $f$ is preserved for unrelated prompts (specificity).
@@ 3.1 Knowledge Editing as a Distribution Mapping Problem
where $f^{*}$ is the edited model, $y^{*}$ is the new desired output (note: not necessarily $o^{*}$), $\mathcal{L}$ a loss assessing the success of the edit by measuring the difference between the edited model’s and the desired output; and $\mathcal{F}$ is the set of all possible edited models. Considering facts are interconnected by a plethora of possible logical implications, it is clear that $\mathcal{X}_{e}$ is quite large, especially when compared to the mere prompt $(s,r)$. This underlines the importance to account for the whole distributions of $\mathcal{X}_{e}$ and $\mathcal{Y}_{e}$ to perform more robust edits.
To approximate the set $\mathcal{X}_{e}$, we aim to generate a set of inputs that are affected by the edit $e$ (i.e. not just $(s,r,o)$ or its paraphrases). It is clear now that our method cannot simply be based on mapping $o\to o^{*}$, because the completions of the inputs in $\mathcal{X}_{e}$ are not limited to the original objects of the edit.
@@ 3.2.2 Mapping and Continuing the Generation
Once the distributions $\mathcal{S}_{e}$ and $\mathcal{T}_{e}$ are collected, we propose to solve Equation 2 using Optimal Transport theory (see e.g. Santambrogio (2015)). Already considered for activation steering in the field of NLP to generate linear counterfactuals Singh et al. (2024); Li et al. (2024b), representation alignment Alqahtani et al. (2021), or more generally steer behavior Rodriguez et al. (2024), optimal transport aims, given two distributions $\mathcal{S}$ (source) and $\mathcal{T}$ (target), at finding a mapping function $m:\mathcal{S}\rightarrow\mathcal{T}$ minimizing some transportation cost, measured traditionally using the Earth Mover’s Distance (EMD) Kantorovich (1960). In particular, and for the sake of simplicity, we focus on linear optimal transport mappings111we use the implementation of the PythonOT library Flamary et al. (2021). That is to say, noting $h$ the last hidden state of the last token of an instance from $\mathcal{S}$, we define the mapping as $m:h\to\mathbf{A}h+\mathbf{b}$, where:
@@ 4.2.1 Logical Implications
Besides the aforementioned baselines, we design another competitor specifically for this experiment. Our objective is to evaluate if our data augmentation strategy described in Section 3.2.1 could also be leveraged to increase the performance of other KE methods. For this purpose, we design CompMEMIT, which leverages MEMIT in order to perform edits that are robust to logical implications. As this method allows for multiple, simultaneous edits, we use the edit prompt and $10$ logical implications prompts generated for SAKE as inputs. Since these prompts all refer to similar knowledge, considering them as simultaneous edits is expected to improve the robustness of the edits while preserving the model’s overall behavior.
The results can be found in Table 1. We observe that, for the two models, SAKE performs better than the weight editing methods and ActAdd along all the dimensions considered. In particular, CompMEMIT performs worse than ROME across all dimensions, suggesting that optimizing over multiple logical implications prompts for the same subject actually negatively impacts the model. On the other hand, while the results of SAKE remain well bellow $100\%$, which may partly be attributed to some inconsistencies we identified in the dataset, they still suggest that the method successfully improves generalization to logical implications.
@@ 4.3 Traditional Editing Evaluation
Finally, we evaluate in this experiment how it performs on traditional KE evaluation metrics. We use the Counterfact dataset Meng et al. (2022a), a commonly used dataset in the KE literature, and calculate the typical KE metrics: accuracy, generality and specificity. We report in Table 3 the scores obtained over the $2000$ first edits. On average, SAKE achieves comparable results to ROME, and outperforms other methods. The most notable improvement is observed on the Generality metric, which may be explained by SAKE’s distribution modeling step, allowing it to account for multiple paraphrase prompts.
@@ 4.4.1 How does SAKE scale with training prompts?
The generation of the in-scope prompts (cf. Sec. 3.2.1), whether agentic or human-based, represents the primary computational cost of performing an edit with SAKE. Hence, an interesting question is how the SAKE’s performance changes as the number of prompts used to collect source and target activations varies, and how many prompts are actually required to achieve satisfactory results. Figure 5 shows the evolution of traditional KE metrics on 150 edits from the Counterfact dataset as a function of the number of prompts $n$ generated in the Distribution Modeling step. We observe that accuracy and paraphrase performance improve as the number of paraphrases increases, since the method becomes more robust to prompts rephrasing. Notably, the specificity score does not decrease with more training prompts, as this number has little impact on the scope detection mechanism. Only 50 paraphrase prompts are thus required to achieve $0.92$ in Accuracy and $0.84$ in Generality.
@@ 6 Limitations
Appendix A Details on SAKE
A.1 Prompts to generate approximation of in-scope input space
@@
A.2 Prompts to collect source and target activations
@@ A.3 SAKE Parameters
    • GPT2-XL and Llama 2-7b, Counterfact: $100$ paraphrases of main edit.
    • GPT2-XL RippleEdits Popular: $100$ paraphrases of main edits, $10$ logical implications augmented with $4$ randomly generated neighbor embeddings for each of the $10$ activations collected.
    • Llama 2-7b RippleEdits Popular: $100$ paraphrases of main edits, around $50$ logical implications with no noise added.
@@ B.2 Competitors
  • CompMEMIT is a competitor designed by us to see if adding logical implications to the main prompt would make the method more robust. In particular, we use MEMIT to edit the model with the main prompt plus $10$ logical implications, then we evaluate it on the appropriate prompts, and then we restore the model to its original form before editing and testing it again.

### v2

@@ Abstract
As Large Langue Models have been shown to memorize real-world facts, the need to update this knowledge in a controlled and efficient manner arises. Designed with these constraints in mind, Knowledge Editing (KE) approaches propose to alter specific facts in pretrained models. However, they have been shown to suffer from several limitations, including their lack of contextual robustness and their failure to generalize to logical implications related to the fact. To overcome these issues, we propose SAKE, a steering activation method that models a fact to be edited as a distribution rather than a single prompt. Leveraging Optimal Transport, SAKE alters the LLM behavior over a whole fact-related distribution, defined as paraphrases and logical implications. Several numerical experiments demonstrate the effectiveness of this method: SAKE is thus able to perform more robust edits than its existing counterparts. The code to reproduce all experiments is made available on a repository [^1] https://github.com/axa-rev-research/knowledge-editing.
@@
Marco Scialanga1,*$\dagger$, Thibault Laugel2,3,*, Vincent Grari2,3, Marcin Detyniecki2,3,4 1EPFL, Lausanne, Switzerland 2AXA, Paris, France, 3TRAIL, LIP6, Sorbonne Université, Paris, France 4 Polish Academy of Science, IBS PAN, Warsaw, Poland *Equal contribution ; Work conducted during an internship at AXA Correspondence: marco.scialanga@epfl.ch, thibault.laugel@axa.com
@@ 2.1 Knowledge Editing
[…]
@@ 3.1 Knowledge Editing as a Distribution Mapping Problem
[…]
@@ 3.2.2 Mapping and Continuing the Generation
Once the distributions $\mathcal{S}_{e}$ and $\mathcal{T}_{e}$ are collected, we propose to solve Equation 2 using Optimal Transport theory (see e.g. Santambrogio (2015)). Already considered for activation steering in the field of NLP to generate linear counterfactuals Singh et al. (2024); Li et al. (2024b), representation alignment Alqahtani et al. (2021), or more generally steer behavior Rodriguez et al. (2024), optimal transport aims, given two distributions $\mathcal{S}$ (source) and $\mathcal{T}$ (target), at finding a mapping function $m:\mathcal{S}\rightarrow\mathcal{T}$ minimizing some transportation cost, measured traditionally using the Earth Mover’s Distance (EMD) Kantorovich (1960). In particular, and for the sake of simplicity, we focus on linear optimal transport mappings222we use the implementation of the PythonOT library Flamary et al. (2021). That is to say, noting $h$ the last hidden state of the last token of an instance from $\mathcal{S}$, we define the mapping as $m:h\to\mathbf{A}h+\mathbf{b}$, where:
@@ 3.2.3 Assessing Edit Scope
Further discussion on the scope detection mechanism is provided in Appendix A.1, as well as a discussion on the overall computational cost of the SAKE (cf. Appendix A.2).
@@ 4.2.1 Logical Implications
Besides the aforementioned baselines, we design another competitor specifically for this experiment. Our objective is to evaluate if our data augmentation strategy described in Section 3.2.1 could also be leveraged to increase the performance of other KE methods. For this purpose, we design CompMEMIT, which leverages MEMIT in order to perform edits that are robust to logical implications. As this method allows for multiple, simultaneous edits, we use the edit prompt and 101010 logical implications prompts generated for SAKE as inputs. Since these prompts all refer to similar knowledge, considering them as simultaneous edits is expected to improve the robustness of the edits while preserving the model’s overall behavior.
[…]
@@ 4.3 Traditional Editing Evaluation
Finally, we evaluate in this experiment how it performs on traditional KE evaluation metrics. We use the Counterfact dataset Meng et al. (2022a), a commonly used dataset in the KE literature, and calculate the typical KE metrics: accuracy, generality and specificity. We report in Table 3 the scores obtained over the 200020002000 first edits. On average, SAKE achieves comparable results to ROME, and outperforms other methods. The most notable improvement is observed on the Generality metric, which may be explained by SAKE’s distribution modeling step, allowing it to account for multiple paraphrase prompts.
@@ 4.4.1 How does SAKE scale with training prompts?
The generation of the in-scope prompts (cf. Sec. 3.2.1), whether agentic or human-based, represents the primary computational cost of performing an edit with SAKE. Hence, an interesting question is how the SAKE’s performance changes as the number of prompts used to collect source and target activations varies, and how many prompts are actually required to achieve satisfactory results. Figure 5 shows the evolution of traditional KE metrics on 150 edits from the Counterfact dataset as a function of the number of prompts $n$ generated in the Distribution Modeling step. We observe that accuracy and paraphrase performance improve as the number of paraphrases increases, since the method becomes more robust to prompts rephrasing. Notably, the specificity score does not decrease with more training prompts, as this number has little impact on the scope detection mechanism. Only 50 paraphrase prompts are thus required to achieve 0.920.920.92 in Accuracy and 0.840.840.84 in Generality.
@@ 6 Limitations
Appendix A Further Details on SAKE: Design and Implementation
A.1 Scope Detection Mechanism
Comparison with a more complex scope detection mechanism.
Relying on a simple distance threshold in embeddings (also used in prior works, e.g.  Hartvigsen et al. (2024)) has several benefits: (1) its simplicity and dependence on a single hyperparameter; (2) its low complexity in terms of calculation (vs. training a dedicated classifier); (3) the flexibility it brings: this way, simultaneous edits are implemented independently, and it becomes very easy to add or remove an edit (vs. needing to retrain a classifier every time); (4) its frugality in terms of training data, especially considering the dimensionality considered. We intended to show that, even with this simple mechanism, good results could still be achieved with our method.
Using a more complex classifier, on the other hand, would likely improve the scope detection, resulting overall in a better generality-specificity tradeoff curve. Especially as the detection is done in high-dimensional embedding spaces: in the experiments conducted, the dimensionality of the space in which the distance is applied is respectively 384, 768, 4096 (prompt model-agnostic embedder, GPT2-XL hidden states, llama2-7b hidden states).
A.2 Computational Cost of SAKE
In this section, we provide a discussion about the computational cost of training and running SAKE.
The following computing characteristics were used for the experiments shown in this paper: MacBook Pro M3 MAX 36GB unified memory, 14-core CPU, 30-core GPU, and 16-core Neural Engine. Training SAKE for an edit relies on the following mechanisms, with associated computation time for GPT2-XL:
  • Generating source and target distributions: this depends heavily on the quickness of inference for the LLM’s API or model considered. In the experiments for this paper, this generally took around 10-20s per edit for 100 training sentences.
  • Collecting activations of these dataset (done on GPU): one edit requires 200 forward passes in total (100 for the source and 100 for the target distributions). This took around 36s for GPT2-XL, 50s for llama2-7b.
  • Learning the Optimal Transport mapping (done on CPU, as the POT library does not provide GPU support): The complexity of learning the OT mapping is $\mathcal{O}(d^{3})$, with $d$ the dimension of the embedding space. For one edit, this step took 0.6s for GPT2-XL, and 12s for llama2-7b.
Comparaison with other methods.
The Optimal Transport mapping leveraged by SAKE requires more computation than other, simpler, techniques, such as ActAdd, which only require computing a difference vector between two prompt representations. ROME on the other hand for instance, requires the computation of a covariance matrix prior to performing the edits, which is also heavy computationally.
A.3 Prompts to generate approximation of in-scope input space
@@
A.4 Prompts to collect source and target activations
@@ A.5 SAKE Parameters
    • GPT2-XL and Llama 2-7b, Counterfact: 100100100 paraphrases of main edit.
    • GPT2-XL RippleEdits Popular: 100100100 paraphrases of main edits, 101010 logical implications augmented with 444 randomly generated neighbor embeddings for each of the 101010 activations collected.
    • Llama 2-7b RippleEdits Popular: 100100100 paraphrases of main edits, around 505050 logical implications with no noise added.
@@ B.2 Competitors
  • CompMEMIT is a competitor designed by us to see if adding logical implications to the main prompt would make the method more robust. In particular, we use MEMIT to edit the model with the main prompt plus 101010 logical implications, then we evaluate it on the appropriate prompts, and then we restore the model to its original form before editing and testing it again.